### v2

@@ Abstract.
We prove rapid mixing for almost all random walks generated by $m$ translations on an arbitrary nilmanifold under mild assumptions on the size of $m$. For several classical classes of nilmanifolds, we show $m=2$ suffices. This provides a partial answer to the question raised in [6] about the prevalence of rapid mixing for random walks on homogeneous spaces.
@@ 1. Introduction
Let $G$ be a simply connected nilpotent Lie group and $\Gamma$ be a co-compact lattice so that $M=G/\Gamma$ is a nilmanifold equipped with the Haar measure $\mu$. A translation on $M$ by $g\in G$ is the map $x\Gamma\mapsto gx\Gamma$. In this paper, we study random walks by a finite set of translations on $M$. These random walks will be defined by an $m\in{\mathbb{N}},\ x\in M$, a set $F:=\{g_{1},\ldots,g_{m}\}\subset G$, and an associated probability vector $\vec{p}=(p_{1},p_{2},\dots,p_{m})$, i.e. $p_{i}>0$ and $\sum p_{i}=1$. The random walk is then a Markov chain where $x_{n}=g_{k}x_{n-1}$ with the probability $p_{k}$. We will let $\langle F\rangle$ denote the semigroup generated by the set $F$. Associated to the set $F$, we have an operator ${\mathcal{L}}:C^{r}(M)\rightarrow C^{r}(M)$ defined by
@@
where $|W|$ is the length of the word $W$ and ${p_{N}(W)=p_{N}(g_{w_{N}}\ldots g_{w_{1}})}=\prod_{i=1}^{N}p_{w_{i}}$ is the probability of having walked by $W$ at time $N$. Since a random walk is exactly defined by its generators, we can view the product $\mu^{m}$ as a measure on the space of all random walks generated by $m$ translations. When we speak of the measure of a set of walks, we mean it with respect to this measure.
@@
A random walk is said to be rapid mixing if given $q\in{\mathbb{N}}$, there are constants $C,r>0$ such that for any $A,B\in C^{r}(M)$ and $N\in{\mathbb{N}}$,
@@
The primary weakness of this theorem is that it sometimes requires more than two generators to guarantee that rapid mixing random walks are of full measure. Within certain special classes of nilmanifolds, we can overcome this deficiency. Among these are the quasi-abelian nilmanifolds, which have previously been studied in the context of parabolic flow dynamics [19, 12]. We also give special attention to triangular nilmanifolds and step-3 nilmanifolds, and ultimately show the following theorem.
@@
It is a classical theorem that for single translations on a nilmanifold, ergodicity, unique ergodicity, and minimality are equivalent and in fact only depend on the irrationality of the abelian projection of the translation. Our result is similar in that we only need an arithmetic property on the restriction of the action in the maximal torus factor. On the other hand, unlike other cases, the arithmetic property we require takes into account how higher-order commutators of the generators will act on $M$. In particular, the argument requires certain polynomials arising from the Lie algebra structure to be linearly independent. In fact, the relevant polynomials only depend on the coordinates of the generators in the abelianization of $G$, and the full measure set of rapid mixing walks we find corresponds to a full measure set of abelian parts that act Diophantinely in some sense (see Definition 3.3). It follows that Theorem A could be strengthened to say that any $m$-tuples in $G$ that project into the aforementioned full measure set of abelian parts generate rapid mixing walks.
A rapid mixing random walk on $M$ also satisfies the Central Limit Theorem (see Appendix B). Several results on central and local limit theorems on nilpotent Lie groups were recently obtained in [15, 2, 3, 5].
We now describe the structure of the paper. Section 2 contains preliminaries on nilpotent Lie groups, Lie algebras, and nilmanifolds. In Section 3, we define $m$-greatness as well as a relevant Diophantine property. We show that the Diophantine property implies rapid mixing (Proposition 3.7). Then, modulo a claim that connects $m$-greatness to the Diophantine property (Proposition 3.4), we show that a full measure set of $m$-tuples on an $m$-great nilmanifold generate rapidly mixing walks. In Section 4, we show Proposition 3.4 by constructing words that act Diophantinely on appropriate tori. In Section 5, we establish that the groups listed in Theorem B are 22-great implying Theorem B. We also show Proposition 5.5 which states that any step $s$-nilmanifold is $s$-great, proving Theorem A. Finally, we provide an example of a Lie algebra that is not $2$-great, showing that our technique is not sufficient to show the suspected optimal result that $N_{G}\!\!=\!\!2$ for all nilpotent groups.
@@ Acknowledgement
The authors appreciate Adam Kanigowski for fruitful discussions and several suggestions for improving the quality of the initial draft. They are grateful to Bassam Fayad and Giovanni Forni for their careful readings and several comments on the draft. M.K. thanks Xuesen Na for his assistance in setting up a computer program for testing calculations.
@@ 2.1. Nilpotent Lie groups and Lie algebras
$\mu(\{\alpha_{01}X_{1}+\ldots\alpha_{0n_{0}}X_{n_{0}}+O(\mathfrak{g}_{1}):(\alpha_{01},\ldots,\alpha_{0n_{0}})\in A\}=\mu(A),$
@@ 2.3. Harmonic Analysis on Nilmanifolds
In this section, we wish to describe the structure of $L^{2}(M)$ for a nilmanifold $M$. To do this, we will need to describe the characters of $M$, denoted by $\widehat{M}$. For any Lie algebra $\mathfrak{a}$, let $\mathfrak{a}^{*}$ represent its dual as a vector space. Since we have a basis for $\mathfrak{g}$ (the Malcev basis), the coordinate functionals $X_{1}^{(1)*},\ldots,X_{n_{s-1}}^{(s-1)*}$ are well-defined and form a basis for $\mathfrak{g}^{*}$. Define $\pi_{p}:\mathfrak{g}\rightarrow{\mathbb{R}}^{n_{p}}$ by $\pi_{p}(V)=(X_{1}^{(p)*}(V),\ldots,X_{n_{p}}^{(p)*}(V))$. Also, let $E_{p}^{*}=\{X_{1}^{(p)*},\ldots,X_{n_{p}}^{(p)*}\}$, and define $\Lambda_{p}$ to be the set of integer linear combinations of elements of $E_{p}^{*}$.
[…]
$H_{\chi}=\{\varphi\in L^{2}(M):\varphi(gx)=\chi(g)\varphi(x)\}.$
We have that
$L^{2}(M)=\bigoplus_{\chi\in\widehat{M}}H_{\chi}.$
Moreover, when $\varphi\in C^{r}(G/\Gamma)$, if we write $\varphi=\sum_{\chi\in\widehat{M}}\varphi_{\chi}$ with $\varphi_{\chi}\in H_{\chi},$ then the weight functions $\varphi_{\chi}$ satisfy $\|\varphi_{\chi}\|_{C^{0}}\leq\frac{\|\varphi\|_{C^{r}}}{\|\chi\|^{r}}.$
@@ 2.4. Main examples
It follows from Definition 2.3 that any quasi-ableian Lie algebra $\mathfrak{g}$ has a basis $(X,Y_{i,j})_{(i,j)\in J}$, satisfying the commutation relations
@@ 2.5. Diophantine Conditions
$|n\cdot v-m|\geq\frac{\gamma}{|n|^{\tau}}.$
In [17], Kleinbock and Margulis show that within submanifolds of ${\mathbb{R}}^{n}$ satisfying a linear independence condition almost every point is in $DC(\gamma,\tau)$} for some $\gamma$ and $\tau.$ In particular, they show the following which is listed as Conjecture $H_{1}$ in their paper.
@@
[17] Let $f_{1},\ldots,f_{n}$ along with the constant function 1 form a linearly independent set of analytic functions from ${\mathbb{R}}^{m}$ to $\mathbb{R}$}. For almost every $x\in{\mathbb{R}}^{m}$, there exist $\gamma,\tau>0$ such that $(f_{1}(x),\ldots,f_{n}(x))\in DC(\gamma,\tau)$.
@@ 3. $m$-greatness and mixing
In this section, we will first define what it means for a nilmanifold to be $m$-great and then show that on these manifolds, almost any $m$-tuple generates a rapidly mixing walk.
@@ 3.1. Linear independence of polynomials
where ${\mathbf{k}}=({\mathbf{k}}_{0},\ldots,{\mathbf{k}}_{p})$, ${\mathbf{k}}_{q}=(k_{q1},\dots,k_{qm})$ and ${\mathbf{k}}_{q}\textbf{V}=\sum_{i=1}^{m}k_{qi}V_{i}.$
The image of $\mathcal{H}_{m,p}$ is exactly the $\mathfrak{g}^{(p)}$ coordinates of the elements $p$-fold brackets of vectors in the span of the $V_{i}$. As we shall see, this will bear relation to the elements $G$ achievable as $p$-fold brackets of elements of the group generated by the $g_{i}$. We use $\mathbf{k}$} instead of $t$ for the first parameter of this map because, while it still makes sense to evaluate this map at any real time, we will be particularly interested in $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)$ that occurs as differences between two different words in the walk. These differences correspond to special values of $\mathbf{k}$}, where $\mathbf{k}$} counts how many times each generator was used at each step of the construction of the words used to obtain the difference $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)$, and thus will always be integral. Note that we have
@@
where each $P_{\ell}({\mathbf{k}},\alpha)$ is a polynomial in the coordinates of $\mathbf{k}$} and $\alpha_{ij}$ (in particular the $\alpha_{0j}$). We will care about when the map $\mathcal{H}_{m,p}$ is degenerate, so in context we have that $\mathcal{D}(\mathcal{H}_{m,p})$ represents the variety
@@ 3.2. $m$-great Lie algebras and mixing
  • A pair of words $W_{1},W_{2}\in\langle F\rangle$ is said to be nice on $G^{(p)}$ if they have equal length and their difference $h:=W_{2}W_{1}^{-1}$ satisfies $h\in G^{(p)}$, and there exist $\gamma,\tau>0$ such that $\pi_{p}(\log(h))\in DC(\gamma,\tau)$.
@@
We postpone the proof to Section 4. We now show that the random walk generated by any $m$-tuple that act Diophantinely on all levels is rapid mixing. The key estimate to show appears in the following theorem:
@@
If a random walk is generated by an $m$-tuple that acts Diophantinely at all levels, then there exist constants $C,a,b>0$ such that for $A\in C^{r}(M)$ with zero mean,
@@
To prove Proposition 3.5, we will show that for any nontrivial character $\chi$, $\mathcal{L}$} has spectral gap on $H_{\chi}$, and, moreover, that the size of the spectral gap has a polynomial lower bound in the norm of $\chi$. We will also use the fact that the size of a $C^{r}$ function’s projection into $H_{\chi}$ is polynomially bounded in the norm of $\chi$. Together, these estimates will allow us to prove Theorem 3.5. We now show the spectral gap.
@@
Assume that $F$ acts Diophantinely at all levels. There exist constants $C_{1},C_{2},\ell,\tau>0$ such that if $\chi$ is a nontrivial character on $M$ and $A\in H_{\chi}$, then
$\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq C_{1}\left(1-C_{2}\left\|\chi\right\|^{-2\tau}\right)^{N/\ell}\|A\|_{C^{0}}.$
@@
Let $A\in H_{\chi}$ where $\chi\in\widehat{M}^{(i)}$. We will now fix words and constants that depend only on $i$ and not on the particular character $\chi$ as follows. Let $W^{i}_{1}$ and $W^{i}_{2}$ be the nice words on $G^{(i)}$ guaranteed by the assumption that the $g_{i}$ act Diophantinely at all levels. Let $\ell_{i}$ be the common length of $W^{i}_{1}$ and $W^{i}_{2}$, and let $\gamma_{i},\tau_{i}$ be the associated Diophantine constants.
@@
Since $A\in H_{\chi}$, we have that
${\mathbf{p}}\left[A(W_{1}x)+A(hW_{1}x)\right]={\mathbf{p}}\left[A(W_{1}x)+\chi(h)A(W_{1}x)\right]=(1+\chi(h))A(W_{1}x).$
Letting $\lambda\in\Lambda_{p}$ to be the nonzero functional such that $\chi=\chi_{\lambda}$, we get that
$|1+\chi(h)|=|1+e^{2\pi i\lambda(\log(h))}|.$
Since by niceness of $W_{1}^{i}$ and $W_{2}^{i}$, $\log(h)\in DC(\gamma,\tau)$, we can apply Lemma A.2 from the appendix, to say that, for some constant $\bar{C}_{i}$ depending only on $\gamma_{i}$, we have that
$|1+\chi(h)|\leq 2-\bar{C}_{i}|\chi|^{-2\tau_{i}}.$
@@
${\mathbf{p}}|A(W_{1}x)+A(hW_{1}x)|={\mathbf{p}}|1+\chi(h)||A(W_{1}x)|\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}}.$
@@
Combining these estimates, we see that on $H_{\chi}$,
$\|{\mathcal{L}}^{\ell_{i}}A\|_{C^{0}}\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}}+(1-2{\mathbf{p}})\|A\|_{C^{0}}\leq(1-{\hat{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}},$
@@
Now we put together the estimates on the behavior of $\mathcal{L}$} on each $H_{\chi}$ and estimate how $\mathcal{L}$} acts on $C^{r}.$
@@
First, we decompose $A$ as $\sum_{\chi\in\widehat{M}}A_{\chi}$ where $A_{\chi}\in H_{\chi}.$ Since $A$ has mean 0, $A_{\chi}=0$ when $\chi$ is the trivial character. Thus,
$\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq\sum_{\|\chi\|^{2\tau}<\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}+\sum_{\|\chi\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}.$
For the first term, we will use the spectral gap of $\mathcal{L}$} from Lemma 3.6. This gives us that when $\|\chi\|^{2\tau}<\sqrt{N}$,
$\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq C_{1}\left(1-\frac{C_{2}}{\sqrt{N}}\right)^{N/\ell}\|A\|_{C^{0}}.$
Thus, since there are only $N^{{\dim(M)/4\tau}}$ characters satisfying $\|\chi\|^{2\tau}<\sqrt{N}$,
$\sum_{\|\chi\|^{2\tau}<\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq C_{1}{N^{\dim(M)/4\tau}}\left(1-\frac{C_{2}}{\sqrt{N}}\right)^{N/\ell}\|A\|_{C^{0}}.$
Setting $C_{3}=C_{2}/\ell$ and $a=\dim(M)/4\tau$, this term is $O(N^{a}e^{-C_{3}\sqrt{N}})\ll N^{{(r-\dim(M))}/{4\tau}}$, so up to changing the constant $K$, we have dealt with the first term. Finally, we observe that since $A\in C^{r}$, the projection $A_{\chi}$ satisfies
$\|A_{\chi}\|_{C^{0}}\leq\|A\|_{C^{r}}\|\chi\|^{-r}.$
Therefore, since $\widehat{M}$ corresponds to a lattice in ${\mathbb{R}}^{\dim(M)}$, we can use the classical bounds on tails of higher dimensional $p$-series to see that for some constant $C_{4}>0$,
$\sum_{\|\chi\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq\sum_{\|\chi\|>N^{1/4\tau}}\frac{\|A\|_{C^{r}}}{\|\chi\|^{r}}\leq C_{4}N^{-(r-\dim(M))/4\tau}\|A\|_{C^{r}}.$
Thus, for the appropriate choice of $K$, $a$, and $b$, the estimate holds. ∎

### v3

@@ Abstract.
We prove rapid mixing for almost all random walks generated by $m$ translations on an arbitrary nilmanifold under mild assumptions on the size of $m$. For several classical nilmanifolds, we show that $m=2$ suffices. This provides a partial answer to the question raised in [6] about the prevalence of rapid mixing for random walks on homogeneous spaces.
Key words and phrases:
Random walks, Rapid mixing, Nilmanifolds, Central Limit Theorem
2020 Mathematics Subject Classification:
22E25, 37A25, 37H05, 60F05, 60G50
@@ 1. Introduction
Let $G$ be a simply connected nilpotent Lie group and $\Gamma$ be a co-compact lattice so that $M=G/\Gamma$ is a nilmanifold equipped with the Haar measure $\mu$. We shall also use $\mu$ to denote the Haar measure on $G$. It will not cause a confusion since the meaning will be always clear from the context. A translation on $M$ by $g\in G$ is the map $x\Gamma\mapsto gx\Gamma$. In this paper, we study random walks by a finite set of translations on $M$. These random walks will be defined by an $m\in{\mathbb{N}},\ x\in M$, a set $F:=\{g_{1},\ldots,g_{m}\}\subset G$, and an associated probability vector $\vec{p}=(p_{1},p_{2},\dots,p_{m})$, i.e. $p_{i}>0$ and $\sum p_{i}=1$. The random walk is a Markov chain where $x_{n}=g_{k}x_{n-1}$ with the probability $p_{k}$. Let $\langle F\rangle$ denote the semigroup generated by the set $F$. Associated to the set $F$, we have an operator ${\mathcal{L}}:C^{r}(M)\rightarrow C^{r}(M)$ defined by
@@
where $|W|$ is the length of the word $W$ and ${p_{N}(W)=p_{N}(g_{w_{N}}\ldots g_{w_{1}})}=\prod_{i=1}^{N}p_{w_{i}}$ is the probability of having walked by $W$ at time $N$.
@@
A random walk is rapid mixing if given $q\in{\mathbb{N}}$, there are constants $C,r>0$ such that for any $A,B\in C^{r}(M)$ and $N\in{\mathbb{N}}$,
@@
Here and below we say that a property holds almost everywhere if it fails on a subset of zero $\mu^{m}$–measure.
[…]
The primary weakness of Theorem A is that it sometimes requires more than two generators to guarantee that rapid mixing random walks are of full measure. Within certain special classes of nilmanifolds, we can overcome this deficiency. Among these are the quasi-abelian nilmanifolds, which have previously been studied in the context of parabolic flow dynamics [19, 12]. We also give special attention to triangular nilmanifolds and step-3 nilmanifolds, and ultimately show the following theorem.
@@
It is a classical theorem that for single translations on a nilmanifold, ergodicity, unique ergodicity, and minimality are equivalent and in fact only depend on the irrationality of the abelian projection of the translation. Our result is similar in that we only need an arithmetic property on associated toral actions. On the other hand, unlike other cases, the arithmetic property we require takes into account how higher-order commutators of the generators will act on $M$. In particular, the argument requires certain polynomials arising from the Lie algebra structure to be linearly independent. However, the relevant polynomials only depend on the projections of the generators in the abelianization of $G$. It follows that Theorem A could be strengthened to say that any $m$-tuples in $G$ that project into the aforementioned full measure set inside the maximal abelian factor generate rapid mixing walks.
A rapid mixing random walk on $M$ also satisfies the Central Limit Theorem (see Appendix B). Several results on central and local limit theorems for a position of a walker on a nilpotent Lie groups were obtained recently in [15, 2, 3, 5]. In contrast, our result concerns additive functions of the walks on nilmanifolds.
We now describe the structure of the paper. Section 2 contains preliminaries on nilpotent Lie groups, Lie algebras, and nilmanifolds. In Section 3, we define $m$-greatness as well as a relevant Diophantine property. We show that the Diophantine property implies rapid mixing (Proposition 3.7). Then, modulo a claim that connects $m$-greatness to the Diophantine property (Proposition 3.4), we show that a full measure set of $m$-tuples on an $m$-great nilmanifold generate rapidly mixing walks. In Section 4, we prove Proposition 3.4 by constructing words that act Diophantinely on appropriate tori. In Section 5, we establish that the groups listed in Theorem B are 22-great implying Theorem B. We also show Proposition 5.5 which states that any step $s$-nilmanifold is $s$-great, proving Theorem A. Finally, we provide an example of a Lie algebra that is not $2$-great, showing that our technique is not sufficient to show the expected optimal result that $N_{G}\!\!=\!\!2$ for all nilpotent groups.
@@ Acknowledgement
The authors thank Adam Kanigowski for fruitful discussions and several suggestions for improving the quality of the initial draft. They are grateful to Bassam Fayad and Giovanni Forni for their careful readings and several comments on the draft. M.K. thanks Xuesen Na for his assistance in setting up a computer program for testing calculations.
@@ 2.1. Nilpotent Lie groups and Lie algebras
$\mu(\{\alpha_{01}X_{1}+\ldots\alpha_{0n_{0}}X_{n_{0}}+\dots:(\alpha_{01},\ldots,\alpha_{0n_{0}})\in A\})={\mathrm{Leb}(A)},$
@@ 2.3. Harmonic Analysis on Nilmanifolds
In this section, we discuss the structure of $L^{2}(M)$ for a nilmanifold $M$. This discussion is based on Kirillov Theory. A general overview of the theory can be found in [4], and [12] includes a similar discussion to ours focused on the theory of filiform nilmanifolds. Let $\mathfrak{a}^{*}$ denote the linear dual to a Lie algebra $\mathfrak{a}\subset\mathfrak{g}$, and let $\mathfrak{a}^{0}$ denote the annihilator of $\mathfrak{a}$ in $\mathfrak{g}^{*}$. For $\lambda\in\mathfrak{g}^{*}$, let $\chi_{\lambda}:G^{(p)}\rightarrow\mathbb{C}$ be the function
$\chi_{\lambda}(g)=e^{2\pi i\lambda(\log(g))}.$
By Kirillov Theory, for every infinite dimensional irreducible unitary representation $H$ of $G$, there is an associated $0\leq p<s$ and $\lambda\in\mathfrak{g}^{(p)*}\cap(\mathfrak{g}^{(p+1)})^{0}$, such that for $\varphi\in{H}$,
$g\varphi=\chi_{\lambda}(g)\varphi\text{ for all }g\in G^{(p)}.$
Observe that the functionals that appear in this description, $\mathfrak{g}^{(p)*}\cap(\mathfrak{g}^{(p+1)})^{0}$, are naturally associated with the functionals in the space $(\mathfrak{g}^{(p)}/\mathfrak{g}^{(p+1)})^{*}\simeq({\mathbb{R}}^{n_{p}})^{*}$.
We need to describe which representations appear in $L^{2}(M)$. Since we have a basis for $\mathfrak{g}$ (the Malcev basis), the coordinate functionals $X_{1}^{(1)*},\ldots,X_{n_{s-1}}^{(s-1)*}$ are well-defined and form a basis for $\mathfrak{g}^{*}$. Define $\pi_{p}:\mathfrak{g}\rightarrow{\mathbb{R}}^{n_{p}}$ by $\pi_{p}(V)=(X_{1}^{(p)*}(V),\ldots,X_{n_{p}}^{(p)*}(V))$. Also, let $E_{p}^{*}=\{X_{1}^{(p)*},\ldots,X_{n_{p}}^{(p)*}\}$, and define $\Lambda_{p}$ to be the set of integer linear combinations of elements of $E_{p}^{*}$. One may note that there is a natural correspondence between the $\Lambda_{p}$ and the characters on the torus $M_{p}$. The significance of $\Lambda_{p}$ is that the $\lambda\in\Lambda_{p}$ are the only functionals in $\mathfrak{g}^{(p)*}\cap(\mathfrak{g}^{(p+1)})^{0}$ with the property that $\chi_{\lambda}$ descends to a well defined function on $M$. Thus, for $\lambda\in\Lambda_{p}$ we can define
$H_{\lambda}=\{\varphi\in L^{2}(M):\varphi(gx)=e^{2\pi i\lambda(\log(g))}\varphi(x)\text{ for all }g\in G^{(p)}\},$
and $L^{2}(M)$ will decompose into a direct sum of such $H_{\lambda}$. In particular, if we let $\Lambda=\Lambda_{0}\cup\bigcup_{p=1}^{s-1}(\Lambda_{p}-\{0\})$, we have that
$L^{2}(M)=\bigoplus_{\lambda\in\Lambda}H_{\lambda}.$
Moreover, when $\varphi\in C^{r}(G/\Gamma)$, if we write $\varphi=\sum_{\lambda\in\Lambda}\varphi_{\lambda}$ with $\varphi_{\lambda}\in H_{\lambda},$ then the weight functions $\varphi_{\lambda}$ satisfy $\|\varphi_{\lambda}\|_{C^{0}}\leq\frac{\|\varphi\|_{C^{r}}}{\|\lambda\|^{r}}.$
@@ 2.4. Main examples
Any quasi-ableian Lie algebra $\mathfrak{g}$ has a basis $(X,Y_{i,j})_{(i,j)\in J}$, satisfying the commutation relations
@@ 2.5. Diophantine Conditions
$|n\cdot v-m|\geq\frac{\gamma}{|n|^{\tau}},$
Let $DC=\bigcup_{\gamma,\tau>0}DC(\gamma,\tau).$
In [17], Kleinbock and Margulis show that within submanifolds of ${\mathbb{R}}^{n}$ satisfying a linear independence condition almost every point is in $DC$. In particular, they show the following which is listed as Conjecture $H_{1}$ in their paper.
@@
[17] Let $f_{1},\ldots,f_{n}$ along with the constant function 1 form a linearly independent set of analytic functions from ${\mathbb{R}}^{m}$ to $\mathbb{R}$}. For almost every $x\in{\mathbb{R}}^{m}$, $(f_{1}(x),\ldots,f_{n}(x))\in DC$.
@@ 3. $m$-greatness and mixing
In this section, we will define $m$-great nilmanifolds and show that on these manifolds, almost any $m$-tuple generates a rapidly mixing walk.
@@ 3.1. Linear independence of polynomials
where ${\mathbf{k}}=({\mathbf{k}}_{0},\ldots,{\mathbf{k}}_{p})$, ${\mathbf{k}}_{q}=(k_{q1},\dots,k_{qm})$ and ${\mathbf{k}}_{q}\textbf{V}=\sum_{i=1}^{m}k_{qi}V_{i}.$ The image of $\mathcal{H}_{m,p}$ is exactly the $\mathfrak{g}^{(p)}$ coordinates of the elements $p$-fold brackets of vectors in the span of the $V_{i}$. As we shall see, this will bear relation to the elements $G$ achievable as $p$-fold brackets of elements of the group generated by the $g_{i}$. We use $\mathbf{k}$} instead of $t$ for the first parameter of this map because, while it still makes sense to evaluate this map at any real value, we will be particularly interested in $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)$ that occurs as differences between two different words in the walk. These differences correspond to special values of $\mathbf{k}$}, where $\mathbf{k}$} counts how many times each generator was used at each step of the construction of the words used to obtain the difference $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)$, and thus will always be integral. Note that we have
@@
where each $P_{\ell}({\mathbf{k}},\alpha)$ is a polynomial in the coordinates of $\mathbf{k}$} and $\alpha_{ij}$ (in particular the $\alpha_{0j}$). We will care about when the map $\mathcal{H}_{m,p}$ is degenerate, since $\mathcal{D}(\mathcal{H}_{m,p})$ represents the variety
@@ 3.2. $m$-great Lie algebras and mixing
  • A pair of words $W_{1},W_{2}\in\langle F\rangle$ is said to be nice on $G^{(p)}$ if they have equal length and their difference $h:=W_{2}W_{1}^{-1}$ satisfies $h\in G^{(p)}$, and $\pi_{p}(\log(h))\in DC$.
@@
We postpone the proof to Section 4. We now show that the random walk generated by any $m$-tuple that act Diophantinely on all levels is rapid mixing. The key estimate appears in the following theorem:
@@
If a random walk is generated by an $m$-tuple acting Diophantinely at all levels, then there exist constants $C,a,b>0$ such that for $A\in C^{r}(M)$ with zero mean,
@@
To prove Proposition 3.5, we will show that for any $\lambda\in\Lambda$, $\mathcal{L}$} has spectral gap on $H_{\lambda}$, and, moreover, that the size of the spectral gap has a polynomial lower bound in the norm of $\lambda$. We will also use the fact that the size of a $C^{r}$ function’s projection into $H_{\lambda}$ is polynomially bounded in the norm of $\lambda$. Together, these estimates will allow us to prove Theorem 3.5. We now show the spectral gap.
@@
Assume that $F$ acts Diophantinely at all levels. There exist constants $C_{1},C_{2},\ell,\tau>0$ such that if $\lambda\in\Lambda$ is a nontrivial functional and $A\in H_{\lambda}$, then
$\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq C_{1}\left(1-C_{2}\left\|\lambda\right\|^{-2\tau}\right)^{N/\ell}\|A\|_{C^{0}}.$
@@
Let $A\in H_{\lambda}$ where $\lambda\in\Lambda_{i}$. We will now fix words and constants that depend only on $i$ and not on the particular $\lambda$ as follows. Let $W^{i}_{1}$ and $W^{i}_{2}$ be the nice words on $G^{(i)}$ guaranteed by the assumption that the $g_{i}$ act Diophantinely at all levels. Let $\ell_{i}$ be the common length of $W^{i}_{1}$ and $W^{i}_{2}$, and let $\gamma_{i},\tau_{i}$ be the associated Diophantine constants.
@@
Since $A\in H_{\lambda}$, we have that
${\mathbf{p}}\left[A(W_{1}x)+A(hW_{1}x)\right]={\mathbf{p}}\left[A(W_{1}x)+\chi_{\lambda}(h)A(W_{1}x)\right]=(1+\chi_{\lambda}(h))A(W_{1}x).$
We then get that
$|1+\chi_{\lambda}(h)|=|1+e^{2\pi i\lambda(\log(h))}|.$
By niceness of $W_{1}^{i}$ and $W_{2}^{i}$, we have that $\log(h)\in DC(\gamma_{i},\tau_{i})$. Since $\lambda$ corresponds to an integral functional, Lemma A.2 from Appendix A says that for some constant $\bar{C}_{i}$ depending only on $\gamma_{i}$, we have that
$|1+\chi_{\lambda}(h)|\leq 2-\bar{C}_{i}\|\lambda\|^{-2\tau_{i}}.$
@@
${\mathbf{p}}|A(W_{1}x)+A(hW_{1}x)|={\mathbf{p}}|1+\chi_{\lambda}(h)||A(W_{1}x)|\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\lambda\|^{-2\tau_{i}})\|A\|_{C^{0}}.$
@@
Combining these estimates, we see that on $H_{\lambda}$,
$\|{\mathcal{L}}^{\ell_{i}}A\|_{C^{0}}\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\lambda\|^{-2\tau_{i}})\|A\|_{C^{0}}+(1-2{\mathbf{p}})\|A\|_{C^{0}}\leq(1-{\hat{C}}_{i}\|\lambda\|^{-2\tau_{i}})\|A\|_{C^{0}},$
@@
Now we put together the estimates on the behavior of $\mathcal{L}$} on each $H_{\lambda}$ and estimate how $\mathcal{L}$} acts on $C^{r}.$
@@
Decompose $A$ as $\sum_{\lambda\in\Lambda}A_{\lambda}$ where $A_{\lambda}\in H_{\lambda}.$ Since $A$ has mean 0, $A_{\lambda}=0$ when $\lambda=0$. Thus,
$\|{\mathcal{L}}^{N}A\|_{C^{0}}\leq\sum_{\|\lambda\|^{2\tau}<\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}+\sum_{\|\lambda\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\lambda}\|_{C^{0}}.$
For the first term, we will use the spectral gap of $\mathcal{L}$} from Lemma 3.6. This gives us that when $\|\lambda\|^{2\tau}<\sqrt{N}$,
$\|{\mathcal{L}}^{N}A_{\lambda}\|_{C^{0}}\leq C_{1}\left(1-\frac{C_{2}}{\sqrt{N}}\right)^{N/\ell}\|A\|_{C^{0}}.$
Thus, since there are at most $N^{{\dim(M)/4\tau}}$ functionals in $\Lambda$ satisfying $\|\lambda\|^{2\tau}<\sqrt{N}$,
$\sum_{\|\lambda\|^{2\tau}<\sqrt{N}}\|{\mathcal{L}}^{N}A_{\lambda}\|_{C^{0}}\leq C_{1}{N^{\dim(M)/4\tau}}\left(1-\frac{C_{2}}{\sqrt{N}}\right)^{N/\ell}\|A\|_{C^{0}}.$
Setting $C_{3}=C_{2}/\ell$ and $a=\dim(M)/4\tau$, this term is $O(N^{a}e^{-C_{3}\sqrt{N}})\ll N^{{(r-\dim(M))}/{4\tau}}$, so up to changing the constant $K$, we have dealt with the first term. Finally, we observe that since $A\in C^{r}$, the projection $A_{\lambda}$ satisfies
$\|A_{\lambda}\|_{C^{0}}\leq\|A\|_{C^{r}}\|\lambda\|^{-r}.$
Therefore, since $\Lambda$ corresponds to a union of lattices in ${\mathbb{R}}^{dim(M)}$ whose dimensions sum to $\dim(M)$, we can use the classical bounds on tails of higher dimensional $p$-series to see that for some constant $C_{4}>0$,
$\sum_{\|\lambda\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\lambda}\|_{C^{0}}\leq\sum_{\|\lambda\|>N^{1/4\tau}}\frac{\|A\|_{C^{r}}}{\|\lambda\|^{r}}\leq C_{4}N^{-(r-\dim(M))/4\tau}\|A\|_{C^{r}}.$
Thus, for the appropriate choice of $K$, $a$, and $b$, the estimate (3.2) holds. ∎